# Mesh-Learner: Texturing Mesh with Spherical Harmonics

### v2

@@ Abstract
In this paper, we present a 3D reconstruction and rendering framework termed Mesh-Learner that is natively compatible with traditional rasterization pipelines. It integrates mesh and spherical harmonic (SH) texture (i.e., texture filled with SH coefficients) into the learning process to learn each mesh’s view-dependent radiance end-to-end. Images are rendered by interpolating surrounding SH Texels at each pixel’s sampling point using a novel interpolation method. Conversely, gradients from each pixel are back-propagated to the related SH Texels in SH textures. Mesh-Learner exploits graphic features of rasterization pipeline (texture sampling, deferred rendering) to render, which makes Mesh-Learner naturally compatible with tools (e.g., Blender) and tasks (e.g., 3D reconstruction, scene rendering, reinforcement learning for robotics) that are based on rasterization pipelines. Our system can train vast, unlimited scenes because we transfer only the SH textures within the frustum to the GPU for training. At other times, the SH textures are stored in CPU RAM, which results in moderate GPU memory usage. The rendering results on interpolation and extrapolation sequences in the Replica and FAST-LIVO2 datasets achieve state-of-the-art performance compared to existing state-of-the-art methods (e.g., 3D Gaussian Splatting and M2-Mapping). To benefit the society, the code will be available at https://github.com/hku-mars/Mesh-Learner.
@@ I INTRODUCTION
However, it suffers from artifacts [11] and fundamental rendering errors (e.g., occlusion error [12]). Furthermore, many applications already have extensive mesh models and require interaction with trained 3DGS or NeRF models. Therefore, there is a need to convert 3DGS or NeRF models into mesh format. However, these conversions not only introduce additional workload but also inevitably lead to losses in rendering quality and geometric integrity. Such limitations lead people to consider whether they can directly learn a texture from the existing mesh, so the learning outcome can be directly applied to a rasterization pipeline without conversion. This is based on the ability to obtain satisfactory meshes with the help of other lidar-based algorithms [13, 14].
@@ II RELATED WORKS
Few works integrate traditional rendering pipelines into the training and rendering process of NeRF or 3DGS. [22] first represented the NeRF model based on textured polygons. They exploit the rasterization pipeline with Z-buffer to achieve pixel-level parallelism, providing each pixel with neural features in texture form. Then, they implement a MLP in the fragment shader to map those features to colors. However, this method is not end-to-end. First, they need to train a NeRF with continuous opacity, then binarize it, and finally bake opacity and features into textures for rendering. In comparison, our method is end-to-end and prioritizes ’high compatibility with traditional pipelines’ as the primary objective. Once the training converges, the output (SH textures) can be imported with the corresponding meshes into a rasterization pipeline for rendering, eliminating the need for any intermediate processing.
@@ III-A SH density
Figure 4: Illustration of world space EWA and the calculation of world space coordinates of SH texels. (a): In the Inside region (red circle), we evaluate texels in the current texture; in the Corner region (blue circle), we evaluate texels in neighbor textures that share the same triangle vertex; in the Edge region (green circle), we evaluate texels in nearby textures that share the same edge. (b): The calculation of world space coordinates for SH texels.
@@ III-D Hybrid SH Interpolation
$u(\mathbf{x})=\begin{cases}\frac{\sum_{i=1}^{N}w_{i}(\mathbf{x})u_{i}}{\sum_{i=1}^{N}w_{i}(\mathbf{x})},&\text{ if }d\left(\mathbf{x},\mathbf{x}_{i}\right)\neq 0\text{ for all }i\\ u_{i},&\text{ if }d\left(\mathbf{x},\mathbf{x}_{i}\right)=0\text{ for some }i\end{cases}$ (1)
@@ III-E World Space EWA Filtering
The EWA’s ellipse coefficients are calculated during the deferred rendering stage; they are originally in local texture space and can be transformed into world space. This is based on the following: 1) we have world space coordinates of the sampled point (i.e., the ellipse center), obtained during the deferred rendering stage. 2) we can calculate the ellipse’s axis lengths in world space by multiplying the texel distance by the SH density. To determine whether the SH texels on neighbouring or self-textures lie within the ellipse in world space, we compute the world coordinates of the SH texels, which were determined in Sec. III-A.
So far, since both the ellipse equation and SH texels’s position are in the same world space, we can determine if a SH texel on the current SH texture or its neighbouring SH textures falls within the ellipse’s range by substituting the coordinates of the SH texel into the ellipse equation. Note that this algorithm assumes that the current mesh and adjacent meshes lie on the same plane. However, this assumption may not always be valid. Therefore, we compare the normals of the current and adjacent meshes; if the angle between the two normals exceeds $15$ degrees, we skip EWA for that adjacent mesh. The strategy for selecting neighbouring texels is similar to that of the hybrid interpolation method: when the SH texture sample point is in the Inside region, we apply EWA solely to the current texture; if in the Edge region, we implement EWA on both the current texture and the neighbouring texture that shares the same edge; and if in the Corner region, EWA is applied to the current texture and the neighbouring textures that share the same vertex.
@@ III-F Adaptive SH Density Strategy
Input: -- Current round PSNR average ${\mathcal{A}_{i}}$, PSNR variance ${\mathcal{V}_{i}}$, and retry times $C_{i}$;
@@ 22             $\mathcal{M}_{i}$ Set Converged
Inspired by [9], we propose an adaptive strategy to address this issue: we assign each mesh a globally low initial SH density and adjust the SH Density changing direction (i.e., increase its density) through a metric method, aiming to adjust each SH density for improved metrics until no further enhancement is achievable. Specifically, after the initial SH density assignment, we train the first round to converge and then evaluate the averages and variances of PSNR based on all input views for each mesh separately. Subsequently, we update SH Densities according to each mesh’s initial density update direction, recalculate SH Texture resolutions, and begin a second round of training. Once it converges, we can obtain another set of metrics (i.e., average and variance of PSNR) of each mesh. We compare these metrics with the previous counterparts; if the average PSNR of a mesh increases or the variance of a mesh’s PSNR decreases, we consider the current update direction to be correct. Otherwise, we consider there’s no need for further increasing the SH Density. Considering the stochastic nature of the training process, we introduce a ”patience” parameter. If the metric does not improve, we allow the mesh another opportunity to optimize in the next training round. Should a mesh’s metrics still not improve under these circumstances, we consider it converged, as it can no longer be optimized to enhance its metrics. In Algorithm 1, $\mathcal{M}_{i}$ refers to a single mesh within the set $\mathcal{M}$ (i.e., all the meshes in a scene), and similarly, $C_{i},\mathcal{A}_{i},\mathcal{V}_{i},\bar{\mathcal{A}}_{i},\bar{\mathcal{V}}_{i},\rho_{i},\bar{\rho_{i}}$ refers to an attribute of a single mesh.
@@ IV-A1 Baselines
Our work is compared to other state-of-the-art methods. For NeRF-based methods, we include InstantNGP [28] and M2-Mapping [23]. For 3DGS-based methods, we select 3D Gaussian Splatting (appearance-oriented) [6] and MonoGS (RGBD-based) [29].
@@ IV-B Replica Dataset
Figure 8: Comparison results on Replica Dataset’s Office-0. The results demonstrate that our work captures more details in both extrapolation and interpolation sequences. In the first row, there’s a zoomed-in view from an extrapolation perspective of a distant clock in the scene. In our results, the small numbers are clear enough to read, while the outcomes of other works are unreadable. In the second row’s zoomed-in view from an extrapolation perspective, our result has sharper edges (blue pad) and more detail (white books). From an interpolation view in the last row, the pattern on the wall in our result is clearer and sharper compared to 3DGS and M2-Mapping.

### v3

@@ Abstract
In this paper, we present a 3D reconstruction and rendering framework termed Mesh-Learner that is natively compatible with traditional rasterization pipelines. It integrates mesh and spherical harmonic (SH) Texture (i.e., texture filled with SH coefficients) into the learning process to learn each mesh’s view-dependent radiance end-to-end. Images are rendered by interpolating surrounding SH Texels at each pixel’s sampling point using a novel interpolation method. Conversely, gradients from each pixel are back-propagated to the related SH Texels in SH Textures. Mesh-Learner exploits graphic features of rasterization pipeline (texture sampling, deferred rendering) to render, which makes Mesh-Learner naturally compatible with tools (e.g., Blender) and tasks (e.g., 3D reconstruction, scene rendering, reinforcement learning for robotics) that are based on rasterization pipelines. Our system can train vast, unlimited scenes because we transfer only the SH Textures within the frustum to the GPU for training. At other times, the SH Textures are stored in CPU RAM, which results in moderate GPU memory usage.
The rendering results on interpolation and extrapolation sequences in the Replica and FAST-LIVO2 datasets achieve state-of-the-art performance compared to existing state-of-the-art methods (e.g., 3D Gaussian Splatting and M2-Mapping). To benefit the society, the code will be available at https://github.com/hku-mars/Mesh-Learner.
@@ I INTRODUCTION
However, it suffers from artifacts [11] and fundamental rendering errors (e.g., occlusion error [12]). Furthermore, many applications already have extensive mesh models and require interaction with trained 3DGS or NeRF models. Therefore, there is a need to convert 3DGS or NeRF models into mesh format. However, these conversions not only introduce additional workload but also inevitably lead to losses in rendering quality and geometric integrity.
Such limitations lead people to consider whether they can directly learn a texture from the existing mesh, so the learning outcome can be directly applied to a rasterization pipeline without conversion. This is based on the ability to obtain satisfactory meshes with the help of other lidar-based algorithms [13, 14].
@@ II RELATED WORKS
Few works integrate traditional rendering pipelines into the training and rendering process of NeRF or 3DGS. [22] first represented the NeRF model based on textured polygons. They exploit the rasterization pipeline with Z-buffer to achieve pixel-level parallelism, providing each pixel with neural features in texture form. Then, they implement a MLP in the fragment shader to map those features to colors. However, this method is not end-to-end. First, they need to train a NeRF with continuous opacity, then binarize it, and finally bake opacity and features into textures for rendering. In comparison, our method is end-to-end and prioritizes “high compatibility with traditional pipelines” as the primary objective. Once the training converges, the output (SH Textures) can be imported with the corresponding meshes into a rasterization pipeline for rendering, eliminating the need for any intermediate processing.
@@ III-A SH Density
Figure 4: Illustration of world space EWA and the calculation of world space coordinates of SH Texels. (a): In the Inside region (red circle), we evaluate texels in the current texture; in the Corner region (blue circle), we evaluate texels in neighbor textures that share the same triangle vertex; in the Edge region (green circle), we evaluate SH Texels in nearby SH Textures that share the same edge. (b): The calculation of world space coordinates for SH Texels.
@@ III-D Hybrid SH Interpolation
[…]
@@ III-E World Space EWA Filtering
The EWA’s ellipse coefficients are calculated during the deferred rendering stage; they are originally in local texture space and can be transformed into world space. This is based on the following: 1) we have world space coordinates of the sampled point (i.e., the ellipse center), obtained during the deferred rendering stage. 2) we can calculate the ellipse’s axis lengths in world space by multiplying the texel distance by the SH Density. To determine whether the SH Texels on neighbouring or self-textures lie within the ellipse in world space, we compute the world coordinates of the SH Texels, which were determined in Sec. III-A. Noted that, we only take adjacent SH Textures into consideration (i.e., those who shares the edge or the vertex with current SH Texture) for moderate computation overhead.
So far, since both the ellipse equation and SH Texels’ positions are in the same world space, we can determine if a SH Texel on the current SH Texture or its neighbouring SH Textures falls within the ellipse’s range by substituting the coordinates of the SH Texel into the ellipse equation. Note that this algorithm assumes that the current mesh and adjacent meshes lie on the same plane. However, this assumption may not always be valid. Therefore, we compare the normals of the current and adjacent meshes; if the angle between the two normals exceeds $15$ degrees, we skip EWA for that adjacent mesh. The strategy for selecting neighbouring texels is similar to that of the hybrid interpolation method: when the SH Texture sample point is in the Inside region, we apply EWA solely to the current texture; if in the Edge region, we implement EWA on both the current texture and the neighbouring texture that shares the same edge; and if in the Corner region, EWA is applied to the current texture and the neighbouring textures that share the same vertex.
@@ III-F Adaptive SH Density Strategy
Input: - Current round PSNR average $\mathcal{A}_{i}$}, PSNR variance $\mathcal{V}_{i}$}, and retry times $C_{i}$;
@@ 22     $\mathcal{M}_{i}$ Set Converged
Inspired by [9], we propose an adaptive strategy to address this issue: we assign each mesh a globally low initial SH Density and adjust the SH Density changing direction (i.e., increase its density) through a metric method, aiming to adjust each SH Density for improved metrics until no further enhancement is achievable. Specifically, after the initial SH Density assignment, we train the first round to converge and then evaluate the averages and variances of PSNR based on all input views for each mesh separately. Subsequently, we update SH Densities according to each mesh’s initial density update direction, recalculate SH Texture resolutions, and begin a second round of training. Once it converges, we can obtain another set of metrics (i.e., average and variance of PSNR) of each mesh. We compare these metrics with the previous counterparts. If the average PSNR of a mesh increases or the variance of a mesh’s PSNR decreases, we consider the current update direction to be correct. Otherwise, we consider there’s no need for further increasing the SH Density. Considering the stochastic nature of the training process, we introduce a “patience” parameter. If the metric does not improve, we allow the mesh another opportunity to optimize in the next training round. Should a mesh’s metrics still not improve under these circumstances, we consider it converged, as it can no longer be optimized to enhance its metrics. In Algorithm 1, $\mathcal{M}_{i}$ refers to a single mesh within the set $\mathcal{M}$ (i.e., all the meshes in a scene), and similarly, $C_{i},\mathcal{A}_{i},\mathcal{V}_{i},\bar{\mathcal{A}}_{i},\bar{\mathcal{V}}_{i},\rho_{i},\bar{\rho_{i}}$ refers to an attribute of a single mesh.
@@ IV-A1 Baselines
Our work is compared to other state-of-the-art methods. For NeRF-based methods, we include InstantNGP [28] and M2-Mapping [23]. For 3DGS-based methods, we select 3D Gaussian Splatting (appearance-oriented) [6] and MonoGS (RGBD-based) [29]. For texture-based methods, we choose the latest work [19] that combines textures with 3DGS.
@@ IV-B Replica Dataset
Figure 8: Comparison results on Replica Dataset’s Office-0. The results demonstrate that our work captures more details in both extrapolation and interpolation sequences. In the first row, there’s a zoomed-in view from an extrapolation perspective of a distant clock in the scene. In our results, the small numbers are clear enough to read, while the outcomes of other works are unreadable. In the second row’s zoomed-in view from an extrapolation perspective, our result has sharper edges (blue pad) and more detail (white books). From an interpolation view in the last row, the pattern on the wall in our result is clearer and sharper compared to 3DGS, M2-Mapping and Textured-GS.